### v1

# Can LLMs Replace Manual Annotation of Software Engineering Artifacts?

Abstract
Experimental evaluations of software engineering innovations, e.g., tools and processes, often include human-subject studies as a component of a multi-pronged strategy to obtain greater generalizability of the findings. However, human-subject studies in our field are challenging, due to the cost and difficulty of finding and employing suitable subjects, ideally, professional programmers with varying degrees of experience. Meanwhile, large language models (LLMs) have recently started to demonstrate human-level performance in several areas. This paper explores the possibility of substituting costly human subjects with much cheaper LLM queries in evaluations of code and code-related artifacts. We study this idea by applying six state-of-the-art LLMs to ten annotation tasks from five datasets created by prior work, such as judging the accuracy of a natural language summary of a method or deciding whether a code change fixes a static analysis warning. Our results show that replacing some human annotation effort with LLMs can produce inter-rater agreements equal or close to human-rater agreement. To help decide when and how to use LLMs in human-subject studies, we propose model-model agreement as a predictor of whether a given task is suitable for LLMs at all, and model confidence as a means to select specific samples where LLMs can safely replace human annotators. Overall, our work is the first step toward mixed human-LLM evaluations in software engineering.
Index Terms:
LLMs, human subjects, evaluation
I Introduction
Given that developer effectiveness heavily depends on suitable tools and processes, researchers are seeking to innovate by improving and automating parts of software engineering. To mention just a few example, there are now many research efforts [1, 2] aimed at code summarization, i.e., techniques that generate a natural language summary for a given piece of code, at detecting bugs and other issues in a program [3], and at determining whether a warning produced by a static analysis tool is actually worth addressing [4].
The usefulness of such tools and techniques ultimately is a question of human cognition. For example, a developer might read an English summary of a method, e.g., in a docstring, and then decide whether and how to use that method. As another example, a developer might look at a warning produced by a static analysis tool, and then decide whether to fix it or not. Because measuring human usefulness is difficult, many research efforts rely on proxy metrics. For the example of code summarization, a summary might rate a reasonable BLEU score, indicating at least some level of similarity to the original, human-written summary. However, even if the BLEU score is good, how can we know if such a summary is clear, precise, relevant to the input code, and human-comprehensible?
To answer such questions, researchers strongly advocate for, and adopt human subject-based evaluations, e.g., when evaluating code summarization techniques [5, 6]. Naturally, this evaluation approach is not just for tools that perform code summarization. Such evaluations are becoming increasingly common in many different settings; indeed in most settings where some aspect of software production is automated using a supporting tool, whether software artifacts are being generated or analyzed, the output from the supporting tool will be used by a human. Since the content & form of the output may affect the ability of the human to properly use the output, it may be sometimes necessary to evaluate the tool using a human-subject study.
Human-subject studies in software engineering are costly. For external validity, such studies require both a representative sample of developers and a representative sample of relevant artifact samples. Each sample may need to be evaluated by multiple humans, to get more stable results. For example, Haque et al. [5], hired professional developers, at \$60 per hour, to rate a total of 420 human and machine-generated code summaries, with three ratings for each sample to gauge inter-rater agreements. Due to the high costs, such evaluations are sometimes done using free or low-cost participants, such as students, which obviously carries some threats, as to whether the results would generalize to professional developers. Yet, even performing a human-subject study with students can be quite time-consuming. In other words, the very reason that motivates researching new tools and techniques in software engineering – the high cost of developers – is challenging the proper evaluation of such tools and techniques.
Given the advances in large language models (LLMs), and their ability to match human performance in a range of tasks [7, 8, 9], the question naturally arises: Can we use LLMs to reduce the cost of human-subject studies in software engineering? Providing a positive answer, even if only for a subset of the effort involved in such studies, could have a huge impact on the practice of evaluations in software engineering. Moreover, improving our understanding of when and how human study participants could be safely replaced by LLMs is an important first step toward mixed human-LLM evaluations.
The above question has arisen in other disciplines, e.g., psychology, linguistics, and medicine [10, 11, 12, 13]. There are several reasons why the software engineering domain is particularly interesting for this exploration, at this time. First, the relative cost: It’s quite costly to recruit representative subjects, whereas LLMs are cheap enough to run hundreds of queries for the cost of using a single human subject for an hour. Second, LLMs are well-trained on software artifacts, both code and code-related natural language, have proven to be useful to a wide range of software engineering tasks, and are now in routine use in the software industry [14, 15, 7, 16]. Third, advances in prompt engineering, such as the impressive abilities of LLMs at in-context learning [17, 18], suggest that with a few illustrative examples, LLMs can execute fairly complex software engineering tasks. Finally, software artifacts are arguably more complex than natural language artifacts: they can require deep knowledge of both application domain and programming; also, they comprise multiple formal & informal elements (e.g, code, summary, warnings).
We study the possibility of substituting human subjects with LLMs when annotating software engineering artifacts. We study ten tasks from five datasets created by prior work, covering research ranging from requirements engineering to reasoning about static analysis warnings. We apply six SOTA LLMs, including both open- and closed-source models, to these tasks, and then compare the LLMs’ responses to those from human subjects. Our findings include:
  • On some tasks, LLMs agree with human subjects about as much as human subjects agree with each other. For example, when judging whether an identifier name and the value it refers to are consistent [19], we observe a mean human-to-human inter-rater agreement of 0.52, and a mean human-to-model agreement of 0.49. This results suggests that LLMs can be a meaningful replacement for humans on some annotation tasks.
  • To help researchers decide whether an LLM may be suitable for a specific annotation task, we investigate the correlation between model-to-model agreement and the ability of LLMs to meaningfully replace humans. We find a strong positive correlation (Spearman correlation of 0.65), which suggests that model-to-model agreement can serve as a relatively cheap-to-obtain proxy to decide whether to involve LLMs in an evaluation.
  • Some examples are easier to annotate than others. For a given example, can we tell whether an LLM can substitute for a human? We find that the LLM’s confidence (its output probability) helps select examples that do not necessarily require human annotators. For example, for the task of judging whether a code snippet and its summary are similar, delegating 50% of the effort performed by one annotator to an LLM (selected based on LLM confidence), does not lead to a statistically significant change of the overall inter-rater agreement.
In summary, we make the following contributions.
  • We are the first to study whether LLMs can replace manual annotation on software engineering artifacts.
  • We present a methodology designed to answer this question and empirical results from applying six LLMs to ten tasks addressed by humans in prior work.
  • We propose a methods for selecting which tasks are suitable for an LLM-augmented evaluation and which samples can be safely delegated to an LLM.
II Background
In software engineering research, evaluations of tools and techniques often rely on qualitative annotations of software artifacts (e.g., code, documentation, bug reports, or newly generated research artifacts [20]). These annotations provide insight into the effectiveness of innovations and their potential impact on developer productivity or software quality. Traditionally, many such evaluations involve human subjects, which requires multiple annotators to ensure consistency and reliability in the data [21].
Inter-rater reliability is vital because it provides a benchmark for the objectivity of the annotations [22]. It is commonly measured using statistical metrics such as Cohen’s $\kappa$ [23] or Krippendorff’s $\alpha$ [24], which measure agreement among annotators by taking into account the possibility of agreement occurring by chance. High values indicate a strong agreement among annotators, suggesting that the annotations are reliable and that the findings are robust.
When annotations are based on a single annotator, there is a significant risk that the findings may reflect personal biases and subjective interpretations, which could lead to inaccuracies and limit the generalizability of the results [25]. In contrast, multiple annotators working independently can achieve a shared understanding, which ensures that the annotations accurately capture the underlying phenomena being studied. However, achieving high inter-rater reliability is effort-intensive and time-consuming, often involving substantial manual work. This challenge is compounded by the variability in human interpretation, which can lead to inconsistencies in the data.
LLMs hold promise to adress these challenges by potentially providing more consistent annotations. Properly trained LLMs could reduce the biases introduced by human annotators and enable the scaling of qualitative analysis to much larger datasets. The integration of LLMs into the annotation process could open up new research opportunities by making it feasible to analyze datasets that were previously too large or complex to annotate manually. In addition, software engineering data is dynamic and constantly evolving, requiring tools that can quickly adapt to new information. LLMs, with their ability to learn from and adapt to new data, could provide up-to-date annotations, complementing the work of human annotators.
To effectively incorporate LLMs into qualitative annotation, it is essential to establish a structured process. This includes deciding when and how to use LLMs, integrating them into existing workflows, distributing tasks between humans and LLMs, and determining the degree of manual work required. This paper addresses these considerations through four research questions, focusing on annotations with pre-existing categories derived from related work or developed in an initial research phase. Future work will explore the potential of LLMs for annotating qualitative data without pre-existing categories (i.e., open coding) within a software engineering context.
III Methodology
In this section, we describe the tasks and datasets, the models under consideration, our research questions, and the methodology to answer these questions. We tried to use the exact same instructions for the models that were used in the original study.
III-A Tasks & Datasets
We select five datasets from previous work, which in total contain ten annotation tasks. The datasets and tasks are selected to represent a diverse range of software engineering research.
Code Summarization Automatic code summarization is an active area aiming to generate helpful summaries of code [3]. One survey-based study reports that 80% of respondents found code comment generation tools useful [26]. 78% agree that these tools help them understand the source code, especially helping code readability in projects with few comments. Several metrics are used to evaluate code summary quality. All metrics aim to indirectly measure human perception of quality, but have limitations. But what criteria are important to evaluate human perception? Haque et al. [5] considered four criteria: accuracy, adequacy, conciseness, and similarity—that should be considered when evaluating code summary quality.
Haque et al. [5] recruited experienced programmers via Upwork, paying them USD60/hr (the applicable market rate). The programmers were presented with 210 functions (from LeClair et al [27]) along with associated human-generated, and model-generated summaries (420 in total). Subjects responded to four questions, related to each of the criteria above. Each participant rated a subset of the 420 samples [^1] https://github.com/similarityMetrics/similarityMetrics, and the authors ensured that each sample was seen by at least three raters.
To evaluate an LLM on this task, we prompt the model with the same guidelines and questions as in the original study. That is, the LLM receives the functions and associated comments, and is tasked to rate its agreement with four statements:
  • Independent of other factors, I feel that the summary is accurate.
  • The summary contains all the important information, and that can help the understanding of the method.
  • The summary contains only necessary information. [^2] As noted in Truong et al [28], LLMs do not handle negation accurately; so this and the above instruction were slightly modified to remove negation, without changing meaning.
  • These two comments are similar.
There are four options: ‘Strongly agree’, ‘Agree’, ‘Disagree’, and ‘Strongly disagree’. To enhance model performance and maintain a specific output format, we use few-shot learning [18, 17]: we prompt the model with a few illustrative query-response pairs and expect the model to respond to the “test” query of our interest. Prompt-size constraints limited us to three shots for most of the experiments.
Name-Value Inconsistencies Variable names are crucial for code understanding [19]. If a variable name does not match the value stored in it, we have an undesirable “name-value inconsistency”.
Patra and Pradel [19] user study had 11 participants, and created a dataset evaluating name-value consistency [^3] https://github.com/sola-st/Nalin. This dataset includes 40 samples rated by all 11 raters. The raters assigned a Likert score from 1 to 5, where 1 indicates difficulty in understanding (or the presence of name-value inconsistency), and 5 indicates ease of understanding. Suppose there is a variable named “name” to which a float, 2.5, is assigned. This is difficult for developers to understand because the expected value was a string. All annotators rated it as difficult to understand. On the contrary, another variable “done”, assigned the boolean value False, is easy to understand and was rated as easy to understand by most annotators. We challenge the LLM to perform the same tasks, again using few-shot prompting.
Causality Causal knowledge supports reasoning about requirements dependencies, particularly for tasks such as test-case construction. Causal relations (e.g., ‘If event 1, then event 2’) are common in system behavior. However, extracting causality from natural language documents is difficult. Fischbach et al. [29] introduced a dataset, built using 6 raters to develop and evaluate a tool for extracting causal dependencies. [^4] https://github.com/fischJan/CiRA The dataset is quite large, with over 10,000 samples. We randomly select 1,000 samples where at least two raters assessed each sentence. The ratings are binary: 1 indicates the presence of a causal relation, and 0 indicates its absence. We then ask the LLM to perform the same annotation task as the original participants based on a few-shot prompt.
Semantic Similarity Functions are often re-implemented several times, with similar functionality, in software projects. However, semantically-similar functions may be implemented in different ways. Finding and merging such functions could reduce maintenance costs. Kamp et al. [30] applied text similarity measures to JavaDoc comments mined from 11 open-source repositories and then manually classified a selection of 857 pairs to create a realistic dataset. [^5] https://github.com/FAU-Inf2/sesame There are three annotation tasks related to functional similarity: goals, operations, and effects. Like for code summarization, each sample was rated by three different people from a pool of eight raters. The original dataset includes some metadata and URLs to the functions but does not contain the function bodies. We retrieve the function bodies and found both bodies for 786 pairs, so we conduct our evaluation on these samples. For each of these pairs, we query LLMs to perform the same three annotation tasks as in the original study.
Static Analysis Warnings Automatic static analysis tools, (e.g. FindBugs), can find coding errors, but do often raise false alarms. Pruning false alarms and presenting only (or mostly) actionable warnings to developers can be beneficial. A dataset by Kang et al[4] provides human annotations on the task of determining whether a particular code change effectively addresses a static analysis warning. The dataset has 1,306 samples that cover different warning categories. [^6] https://github.com/soarsmu/SA_retrospective/tree/main Two annotators assigned each sample one of three labels: open, closed, or unknown. A static analysis warning is considered “closed” if it was reported in a previous software revision but not in a subsequent reference revision, indicating that the problematic code was altered or removed. Conversely, a warning is labeled as “open” if it appears in both the current testing revision and a later reference revision, suggesting that the warning was not actionable or it was ignored by the developers. Lastly, a warning is marked as “unknown” if the file containing the warning was deleted or modified in a (possibly unrelated) way in the reference revision, making it difficult to confirm whether the warning was actionable. Given the extensive metadata, the prompts for this problem are the longest. Due to rate limitations imposed by some models (e.g., Claude and Gemini) and the need to manage costs, we use 200 samples, chosen uniformly at random, for our experiments. For each of these samples, we ask LLMs to perform the annotation task.
Detailed prompts (including guidelines and few-shot samples) for each dataset are provided in the supplementary material.
III-B Models under Consideration
We use both closed and open models. Among closed models, we chose GPT-4, Claude-3.5-Sonnet, and Gemini-1.5-Pro, which are all the best models from their respective families or organizations at the time of the work. We also used the GPT-3.5 model. From the open models, we chose Llama3 (70B) and Mixtral (8x22B), which are also the latest models from the open-model family.
III-C Research Questions
Our goal is to study whether LLM prompting can subsitute for human annotation labour, for artifact annotation tasks in SE. Specifically, we want to understand how annotator (inter-rater) agreement changes when substituting human annotations with model annotations. In the first research question, we examine all three categories of inter-rater agreement: human-human, human-model, and model-model. Specifically, if human-model inter-rater agreement is similar to that of human-human agreement, it may be possible to substitute (some) human ratings with model ratings. We will also examine the inter-model agreement and how it changes across datasets.
RQ 1. When humans are replaced by LLMs to provide answers for “human-rater” questions in software engineering research, what level of agreement is observed between humans and models?
In Section IV-A we discuss the findings of human-human, human-model and model-model inter-rater agreement. Depending on the dataset and annotationt task, we find varying levels of agreement. This raises the question of how one might decide whether a particular task is amenable to replacing human effort with an LLM.
RQ 2. How can we determine if LLMs are NOT usable for a specific task?
In contrast to RQ2, where we consider basic feasibility of replacing human effort with a model, the next question focuses on identifying specific samples where we can replace one human rating with a model rating. Although a model may not perform well for all samples, selecting samples where models are successful can still significantly reduce human effort.
RQ 3. How can we determine if an answer from an LLM for a specific sample is likely to be in agreement with a human answer?
In case we have determined if and when to delegate parts of a human-subject study to LLMs, the final research question is about the cost benefits of doing so.
RQ 4. How much human annotation effort could be saved without sacrificing inter-rater agreement?
III-D Evaluation Methodology
For RQ1: With each task, we record a) number of annotators and b) samples rated by each annotator. Note that an annotator may not have annotated all the samples in the dataset but rather a subset. [^7] For example, Haque et al used six annotators, but each sample only received three annotations. After identifying the samples annotated by each human, we compute the inter-rater agreement metric Krippendorff’s $\alpha$ for each human-human, human-model, and model-model pair. To clarify, if a human (P1) has rated 5 samples {s1, s2, s3, s4, and s5}, and another human (P2) has rated 5 samples {s3, s4, s5, s6, s7}, we consider only {s3, s4, s5} for computing Krippendorff’s $\alpha$ for P1 and P2. There are other inter-rater agreement metrics, but Krippendorff’s $\alpha$ is applicable to any number of annotators, each assigning one value to one unit of analysis, to incomplete (missing) data, to any number of values available for annotating a variable, and to different kinds of annotations (binary, nominal, ordinal, interval, ratio, etc.). Given the diversity of the datasets we study, this metric allows us to uniformly apply inter-rater agreement to all datasets.
For RQ2: As mentioned earlier, we compute human-human, human-model, and model-model agreement $\alpha$. Note that model-model agreement is cheap and automated to determine. To answer this question, we examine the model-model agreements and human-model agreement. If these two are correlated, then model-model agreement can help judge the possibility of having high human-model agreement.
For RQ3: To determine whether asking an LLM for a specific sample is a good idea, we use the output probability of a model’s answer as a proxy for the model’s confidence. We then study the impact of replacing those human answers where an LLM gives the highest-confidence responses with an LLM answer. To assess the impact, we investigate how the inter-rater agreement changes if we replace one human annotator with an LLM for a specific fraction of all samples.
For RQ4: We measure the human effort one could save without reducing inter-rater agreement in a statistically significant way. Our unit of effort here is the work required, on average, to label one sample. Saving 100% of the effort for one rating corresponds to getting one annotation on each sample in a dataset from an LLM. Note that this saving typically saves at least one, and sometimes multiple human participants. For example, if a problem requires a total of 300 annotations for 100 samples, then saving 100% rating effort means to replace 100 annotations (which otherwise would be annotated by one or more humans).
IV Results
IV-A RQ1: Level of Agreement between Humans and Models
We now present the observed level of agreement between humans and models across different datasets.
Datasets
Human to Human
Inter-rater agreement
Human to Model
Inter-rater agreement
Model to Model
Inter-rater agreement
Mean Median Mean Median Mean Median
Accuracy (Code Summarization) 0.38 0.44 0.48 0.48 0.76 0.78
Adequacy (Code Summarization) 0.40 0.39 0.41 0.38 0.74 0.73
Conciseness (Code Summarization) 0.24 0.26 0.21 0.26 0.74 0.75
Similarity (Code Summarization) 0.64 0.66 0.44 0.43 0.68 0.68
Name-Value Inconsistencies 0.52 0.52 0.49 0.48 0.66 0.67
Causality 0.44 0.49 0.22 0.28 0.39 0.25
Goals (Semantic Similarity) 0.83 0.83 0.77 0.78 0.82 0.81
Operations (Semantic Similarity) 0.74 0.77 0.67 0.68 0.77 0.79
Effects (Semantic Similarity) 0.71 0.73 0.64 0.65 0.69 0.64
Static Analysis Warning 0.80 0.80 0.15 0.23 0.12 0.09
TABLE I: Summary of Inter-Rater Agreements. Three Best Performing Models (i.e., Claude, Gemini and GPT-4) have been Used to Report the Agreement in this Table.
Code Summarization Figure 1 shows heatmaps of inter-rater agreement over individual ratings on experimental samples, for human-human, human-model, and model-model pairs for code summarization criteria: Accuracy, Adequacy, Conciseness, and Similarity. The original study used six human raters. We add ratings from six different models, along with (for comparison) an automated rater which randomly assigned labels to the samples. Figure 1-(a) shows three zones representing three categories of inter-rater agreement, based on available data. [^8] Some human raters did not rate the same samples.
Model-model agreement is high, for all criteria, especially for the three large models (GPT-4, Gemini, and Claude). Table I indicates that the mean Krippendorff’s $\alpha$ is 0.68-0.76. Second, we see that human-model and human-human agreements are in similar ranges, 0.24-0.40 and 0.21-0.48 for the first three categories. Because of the way models are trained, we can expect powerful models to tend to reflect the “majority opinion” they learn from the training corpus.
For similarity, human-human agreement is higher than human-model agreement. Although smaller models are also useful, they show some bias towards specific answers. For example, Llama-3 consistently chose between “Strongly agree” and “Strongly disagree”, ignoring the options “Agree” and “Disagree”. Similarly, human rater “P2” was not particularly satisfied with the conciseness of the program, resulting in lower inter-rater agreement with all humans and models.
To summarize, for all criteria, we observed generally similar human-human and human-model inter-rater agreements, with a few exceptions.
(a) Accuracy
(b) Similarity
Figure 1: Inter-Rater Agreement (Krippendorff’s Alpha) for Code Summarization Accuracy and Similarity. For Adequacy and Conciseness, we have similar plots to Accuracy; thus, we omitted them from the paper due to page constraints.
Name-Value Inconsistencies We find that human-model and human-human agreement are quite similar for this task (mean Krippendorff’s $\alpha$ 0.49 vs. 0.52). Figure 2 shows a heat-map for all 11 humans, 6 models, and a random rater. We have very high values in all three zones: human-human, human-model, and model-model. Here, too, we find higher agreement between the top 3 models (mean Krippendorff’s $\alpha$ 0.66). Note that we will consider only the top three models for reporting Krippendorff’s $\alpha$ from here on because we found that only these three models do not show bias or preference towards any specific answer.
Figure 2: Inter-Rater Agreement for Name-Value Inconsistencies.
Causality The models struggle more to detect causality, relative to the two prior tasks (Figure 3). The mean human-model agreement (0.22) is much less than human-human (0.44). The mean model-model agreement (0.39) is less than 0.5. To summarize, we found lower level of agreement between human-model and model-model for this task.
Figure 3: Inter-Rater Agreement for Causality.
Semantic Similarity For the functional semantic similarity dataset in all categories (e.g., goals, operations, and effects), we have observed high agreement for all human-human (0.71-0.83), human-model (0.64-0.77), and model-model (0.69-0.83) pairs (see Table I and Figure 4). We found the highest level of inter-rater agreement in all three pairings.
Figure 4: Inter-Rater Agreement for Semantic Similarity Evaluation Criteria ‘Goals’. For other criteria (‘Operations’ and ‘Effects’), we have very similar heatmaps; thus, they were omitted from paper.
Static Analysis Warning For the static analysis warnings rating task, we have only two human raters, and they are in strong agreement (Krippendorff’s $\alpha$ 0.80). However, human-model (0.15) and model-model (0.12) agreements are low (Figure 5). These findings suggest that LLMs cannot safely substitute for human ratings in this task.
Figure 5: Inter-Rater Agreement for Static Analysis Warning.
We note that human-model agreement varies from task to task. While it resembles human-human agreement values for some tasks (code summarization, name-value inconsistencies, semantic similarity), it is lower for other tasks (causal relation detection and static analysis warnings).
IV-B RQ2: LLMs’ Applicability for a Specific Task
We observed that for a majority of tasks, human-model inter-rater agreement is very similar to human-human agreement, while for some tasks, it is lower. What about model-model inter-rater agreement, which does not require any human effort? Figure 6 shows that the mean model-model agreement of the top 3 models is positively correlated with the mean human-model agreement, with a Spearman correlation of 0.65 and a $p-value<0.05$. This suggests that higher model-model agreement is helpful in indicating good human-model agreement; this observation is potentially valuable in deciding whether to replace humans with models. There is an outlier (high model-model agreement but low human-model agreement) in the bottom-right corner, but for this case (code summarization conciseness), human-human agreement is also low (0.24) suggesting that human ratings for this task per se are not consistent. Our results suggest that if multiple LLMs reach similar solutions independently, then LLMs are likely suitable for the annotation task.
Figure 6: Human-Model Inter-Rater Agreements is Positively Co-Related with Model-Model Inter-Rater Agreements with $p-value<0.05$. Human-Human agreement are shown in blue text.
IV-C RQ3: LLMs’ Applicability for a Specific Sample
If we observe the tasks where the models and humans are reasonably agreeing with each other, the model-model agreements are quite high (0.66-0.82). For two datasets, causality detection and static analysis warning, the inter-model agreements are too low (0.39 and 0.12). Now we will see whether we can replace one human rating in each sample with a model. We will discuss the results with respect to the GPT-4 model output from here on, because this model allows us to observe the output probability and is the best-performing model overall.
Figure 7(a & b) shows how inter-rater agreement changes with increasing fraction of samples from GPT-4 for code summarization accuracy and similarity. For adequacy and conciseness, we have similar plots to accuracy, so we have omitted them from the paper due to page constraints. Consider Figure 7(a), where we gradually replace the human ratings with model ratings. On the x-axis, 10% means we take 10% of the samples and replace one randomly chosen rating with GPT-4 output, then observe the inter-rater agreement (now including 10% model ratings).
We can choose the 10% samples in two ways: based on GPT-4 output probability or randomly. The red lines show the results when we choose the samples based on probability, and the blue line indicate when samples are chosen randomly. After selecting the samples, we have three ratings for each sample. We randomly select a position, replace the rating with GPT-4 output, and calculate Krippendorff’s $\alpha$ as in RQ1. We repeat the process 100 times for each selection criterion. For each data point on the x-axis, we present the mean value and 95% confidence interval. For human-human agreement, we present the inter-rater agreement with a dotted line and have drawn the confidence interval using bootstrapping by randomly selecting 50% of the samples and repeating the process 1,000 times.
(a) Accuracy
(b) Similarity
Figure 7: Inter-Rater Agreement with increasing % of Samples from GPT-4 for Code Summarization Accuracy and Similarity. The dotted line indicates the inter-human agreement for that task.
In general, the probability-based selection criterion works better than randomly selected samples. However, at 100%, there is no difference between them because all the samples have been considered for replacement. For the accuracy task of code summarization (Figure 7), we find that the inter-rater agreement does not change at all, even when we replace one rating in all the samples. It slightly increases with the probability-based selection approach, the improvement is insignificant and within the 95% confidence interval of human-human agreement. For the similarity task (Figure 7-b), we have seen that GPT-4 maintains inter-rater agreement up to 50% of the samples while chosen based on probability, potentially saving 16.5% of overall human effort.
Figure 8: Inter-Rater Agreement with increasing % of Samples from GPT-4 for Name-Value Inconsistencies.
For name-value inconsistencies (Figure 8) and functional similarity (figure omitted for space constraints), we observed that we can replace one rating in all samples with GPT-4 and still maintain the same inter-rater agreement. Note that for name-value inconsistency, we have few samples. Therefore, the confidence interval of human-human agreement is much wider.
Figure 9: Inter-Rater Agreement with increasing % of Samples from GPT-4 for Causality.
Figure 10: Inter-Rater Agreement with increasing % of Samples from GPT-4 for Static Analysis Warning.
For causality and static analysis warning, we have low inter-model agreement (0.39 and 0.12). We have already discussed that we should expect low human-model agreement because they are correlated. Figures 9 & 10 show that the inter-rater agreement decreases as we increase the samples from the GPT-4 model, indicating that the model may not be applicable in this setup. However, using the confidence-based approach, we can still see that GPT-4 maintains similar inter-rater agreement, up to 60% for the causality dataset and 50% for the static analysis dataset, showing some promise for using GPT-4 for partial datasets. The confidence interval for these datasets up to that point completely overlaps with the confidence interval of the human-human rating. High-probability samples are likely to be more correct or align better with human preferences compared to other samples.
Datasets
#of Rating for
Each Sample
% Effort Saved
for One Rating
% Effort Saved
for Overall Process
Accuracy (Code Summarization) 3 100% 33%
Adequacy (Code Summarization) 3 100% 33%
Conciseness (Code Summarization) 3 100% 33%
Similarity (Code Summarization) 3 50% 16.5%
Nalin 11 100% 9%
Causality 2 60% 30%
Goals (Sesame) 3 100% 33%
Operations (Sesame) 3 100% 33%
Effects (Sesame) 3 100% 33%
Static Analysis Warning 2 50% 25%
TABLE II: % of Human-Effort Saved for One Rating and Overall Annotation Process.
IV-D RQ4: Human Effort to Potentially Save
By replacing one human rating with an LLM’s answer for some fraction of all samples, as studied in RQ3, we can save some human effort. The saved effort depends on two factors: First, the number of ratings required per sample. As we replace only one human’s rating for each sample, more required ratings per sample means less potential for saving human effort. Second, the fraction of samples for which we replace a human rating with an LLM’s rating. To decide on this fraction, we use the results from RQ3 to determine the maximum fraction that keeps the resulting inter-rater agreement statistically indistinguishable from a human-only study. For example, in Figure 7-b, asking an LLM for help for up to 50% of all samples keeps the agreement within the zone marked with gray background, but going beyond that fraction would change the inter-rater agreement in a statistically significant way.
Based on this reasoning, Table II summarizes the amount of effort that could potentially be saved. For the accuracy task of code summarization, we can save 100% of the effort to obtain one rating per sample, which is 33% of the overall rating effort for this tasks. For name-value inconsistencies, we can save only 9% of the overall effort because there are 11 ratings per sample, whereas for the semantic similarity dataset, we can save up to 33% of the effort because it involves only three ratings per sample. Note that we ignore the few-shot labeling effort (which is 3-4 samples only) for this estimation in Table II. Overall, for seven of the ten tasks, we can safely replace one human rater with a model.
V Discussion & Limitations
Figure 11: Steps to Decide Human Rating Replaceability.
Deciding Whether and When to Ask an LLM Our findings suggest that we can replace one human rating by a model for a significant number of samples (50%-100%). However, it appears risky to fully replace all humans, even though the probability of the model output at least somewhat indicates the quality of annotations. Based on our findings, we propose the process as given in Figure 11, which involves two steps: 1) Create few-shot examples (3–4 in our setup) to query multiple strong LLMs with all samples, and compute the model-model agreement. 2) If the agreement is high ($>0.5$ based on our data), one can safely replace one human rating per sample with an LLM-provided answer. Otherwise, selectively replace one rating only for those samples where the LLM gives a high-confidence answer.
Beyond saving human effort, our findings can also help extend a dataset by automatically labeling additional samples. We should emphasize that replacing more than one human can inflate inter-rater agreement because model-model agreements are much higher. This may not fulfill or reflect the original goals of the data annotations.
Confidence Distribution over Samples? We did find that picking the samples based on model confidence yields better results, even when model-model agreement is low, for almost 50% of the data. However, the question remains: what should be our cutoff confidence? Figure 12, which orders fraction of the sample, at decreasing confidence levels, shows that GPT-4 is usually quite confident with its solutions. However, after a certain point, the probability starts to decrease. For causality and static analysis warning, we can see that for the first 50%-60% of samples, the model maintains a high output probability, around 0.80. Thus, the probability cutoff point can be computed when we have low model-model agreement.
Figure 12: Change of Output Probability (by GPT-4) with % of Samples Sorted Based on GPT-4 Model’s Output Probability. Some datasets were omitted for clarity of the plot.
Deciding Whether All Humans Are Replaceable In RQ3, we asked if one rating in all samples can be replaced by an LLM. Can we replace all humans, but just for specific samples? First, we clarify when we can replace all the humans for a sample. Note that there are very few samples where all annotators would agree. For name-value inconsistencies, no sample exists where all annotators agree on a particular rating. Therefore, we will use majority voting to decide whether all humans are replaceable for a sample or not. For a dataset with 2 annotators, the model needs to agree with both annotators, and for a dataset with 11 annotators, the model needs to agree with at least 6 annotators to replace all humans.
Figure 13: Change of Fraction of Samples where GPT-4 Agress with Majority Voting for % of Samples Sorted Based on GPT-4 Model’s Output Probability.
Figure 13 shows that samples with higher GPT-4 output probability are more likely to agree with the majority of annotations; however, even with 10% of the samples, there is some possibility of error. Apart from semantic similarity, across all datasets, we can’t find any probability split where the model consistently agrees with majority voting. Therefore, we conclude that we cannot replace all the humans for a sample with the current models. We can replace humans for selective samples, but there is some risk associated with it. More study is required for this; we leave it for future research.
Limitations We consider only 10 tasks datasets and 6 models. It is quite difficult to find publicly available datasets where the ratings from individual annotators are still available; many artifacts contain only the final annotation after resolving disagreements. Even though we see pretty consistent results across the datasets and models, finding may vary with others. Only Krippendorff’s $\alpha$ was applicable to all studied datasets. Cohen’s $\kappa$ is another popular measure but is applicable to only two raters. Using few-shot learning means that some human effort is required initially. However, labeling 3–4 samples is much cheaper compared to labeling all the samples.
Further Investigations It is now well-established that few-shot learning performs better than zero-shot learning. We also tried zero-shot learning in our preliminary stage and found that the human-model agreement was not satisfactory, even for datasets where few-shot learning performed very well. We tried repeated sampling (at higher temperature): the models mostly generate the same samples. We could improve the multi-sample technique by using chain-of-thought [31] or self-consistency [32], but we do not have the original thoughts used by the original human annotators. In future studies, we strongly recommend collecting human thoughts, which could be used for few-shotting.
VI Related Work
ML and LLMs in Software Engineering Learning-based approaches and LLMs have been applied to various software development tasks [33], including code completion [34, 35, 36], test generation [37, 38, 39, 40], fuzzing [41, 42], automated repair [43, 44, 45, 46, 47] and coding agents [48, 49, 50, 51], and type prediction [52, 53, 54, 55]. We follow this trend, but deviate by not targeting a specific task, but by studying the potential for automating human-subject (manual) evaluations, using LLMs.
LLMs to Complement or Replace Human Annotators Recent and concurrent work explores the idea of complementing or replacing human annotators with LLMs in domains beyond software engineering. Work in natural language processing (NLP) compares LLMs and human participants, typically recruited via crowd-working platforms, and report mixed results. While some studies show that adding LLM labels can improve the aggregated labels of a dataset [56] and sometimes outperform crowd-workers [57], others note that LLMs fail to accurately represent differences between demographic groups [58]. Others suggest to use LLMs to perform sub-tasks in a crowd-sourcing pipeline [59]. Bavaresco et al. [13] propose a benchmark of NLP datasets to compare human and LLM performance. We envision our work to serve as such a benchmark in software engineering. To the best of our knowledge, our paper is the first to carefully study LLMs for annotating software engineering artifacts, and propose actionable guidelines for when and how to use LLMs to create annotations.
Collaborative Human-LLM Annotation Our suggested workflow (Section V) relates to work on human-LLM collaboration for data labeling [60]. Some papers [61, 12] propose to first query an LLM, and then let humans refine some of the LLM-provided labels. Instead, we find that inter-model agreement and LLM output probabilities provide an effective way of deciding which labeling tasks can be safely delegated to a model.
LLMs to Automate Research Tasks Our work fits into a broader stream of work on partially automating particular sub-tasks of performing academic research. Liang et al. [62] study to what extent LLMs can replicate empirical software engineering research by determining assumptions made in a study and by writing data processing code. While we share the idea of using LLMs to automate a research task, we focus on a different kind of task. Other work proposes to let LLMs generate open-ended questionnaire responses for HCI-style evaluations [63], and shows a set of LLMs can reproduce human-study results in economics, psycholinguistics, and social psychology [10]. We expect to see more work on delegating sub-tasks of conducting research to LLMs, e.g., for performing human-intensive evaluations and editing scientific texts.
LLMs for Data Synthesis and Assessment Beyond using LLMs to partially automate evaluations that would otherwise be performed by humans only, others propose to use LLMs to synthesize additional data to train or fine-tune models for text annotation tasks [64, 65]. There is a recent related survey [66]. A key difference from our work is that training and fine-tuning datasets must be large-scale, but some amount of noise is acceptable, whereas annotations in human-subject studies are typically of smaller scale, but should have high confidence. LLMs have also been proposed as judges of outputs generated by other LLMs, either with a single LLM as the judge [67] or with a set of LLMs that discuss until reaching an agreement [68]. These efforts are primarily about determining human preference, i.e., tasks where different humans may legitimately disagree, whereas many annotation tasks in software engineering have an objectively correct answer that humans can eventually agree upon.
VII Conclusion
In this paper, we investigate if LLM responses can substitute for human raters in software engineering annotation tasks, and how such substitutions may affect inter-rater agreement. We find that human-model agreements can be fairly consistent with human-agreements, for the majority of the settings we studied. We also find model-model agreement is a good indication of human-model agreement, suggesting that when powerful models trained on giant human corpora agree with each other, they tend to agree also with humans. Finally, we find that an LLM’s confidence (output probability) for a given sample output is a good indication of whether LLM output agrees with majority human rating, for that sample. We caution that the paper considers only discrete (multiple choice) LLM responses, not free-form annotations; furthermore, we have not studied issues of model output bias, nor issues of demographics. Finally, our scripts and datasets are made available anonymously at https://zenodo.org/doi/10.5281/zenodo.13146386